OpenStudy (blackstreet23):

What was the calculus formula to determine the volume of a solid? I don't mean solid of revolution. Just regular solid.

OpenStudy (blackstreet23):

volumes of cross sections

OpenStudy (freckles):

Are you sure you aren't looking for: \[\text{ Volume }=\int\limits_a^bA(x) dx \text{ where } A(x) \text{ is the area of the cross section } ?\]

OpenStudy (blackstreet23):

yes that is it ! thank you !

OpenStudy (freckles):

that's also the formula used in solids formed by revolution of some curve about some axis

OpenStudy (lynfran):

u forget to square it..i think @freckles

OpenStudy (freckles):

you know where we take circles as the cross section

OpenStudy (freckles):

no A(x) is the area you don't want to square the area

OpenStudy (lynfran):

o ok

OpenStudy (freckles):

if the cross section is a circle then A(x)=pi*r^2

OpenStudy (freckles):

where r is a function of x

OpenStudy (freckles):

though we could look at things in terms of y too not just x

OpenStudy (freckles):

you know depending on what suits us better

OpenStudy (blackstreet23):

but how is that the formula for solids of revolution? I mean the formula for solids of revolution is \[\int\limits_{a}^{b}(\pi*f(x)^2) dx\]

OpenStudy (lynfran):

i think the f(x) is actually the radius

OpenStudy (blackstreet23):

f(x) is the function

OpenStudy (blackstreet23):

i mean the height of the funtion

OpenStudy (lynfran):

yes which is intern the radius so A=pi*r^2

OpenStudy (lynfran):

and freckles just use the Area instead...i think

OpenStudy (blackstreet23):

I guess i am kind of confuse lol

OpenStudy (lynfran):

ok lets wait on till @freckles come online...

OpenStudy (freckles):

what's the question?

OpenStudy (freckles):

area of a circle is what I called A(x)

OpenStudy (freckles):

or the are of the cross section is A(x)

OpenStudy (freckles):

if the cross section is a circle then A(x) is pi*r^2 where r is the radius f(x) is the distance from the curve to the x-axis so f(x) is the radius if the axis of rotation is the x-axis

OpenStudy (freckles):

 Say we want to take this curve on [a,b] and rotate it about the x-axis